OpenStudy (artfis):

I'm doing some studying and I need some step by step directions for this one. Solve for i. A=P(1+i/100)^2

OpenStudy (anonymous):

1st step: Divide both sides by P 2nd step: Square root both sides

OpenStudy (anonymous):

A/P = (1 + i / 100)^2 sqrt(A/P) = 1 + i / 100 100 * sqrt(A/P) = 1 + i 100 * sqrt(A/P) - 1 = i \[(100 \times \sqrt{A/P}) - 1\] That's it if i'm not mistaken

OpenStudy (anonymous):

i = that last line...

OpenStudy (anonymous):

Think you messed up on line 3....it's 1 + i/100...not (1 + i)/100 So when you multiply both sides by 100 line 3 right side of the = should be 100 + i

OpenStudy (ranga):

Going from line 2 to line 3, subtract 1 first and then multiply by 100.

OpenStudy (anonymous):

Exactly, if it was meant to be 1 + (i/100) then that's correct. I assumed that it was (1 + i)/100

OpenStudy (anonymous):

It's the formula for compound interest applied for 2 periods 100 * sqrt(A/P) = 100 + i 100* sqrt(A/P) - 100 = i or i = 100(sqrt(A/P) - 1) factoring out the 100

OpenStudy (artfis):

Ok. Thank you everyone. Much appreciated.

OpenStudy (anonymous):

Anytime, Good luck with your studies. Thanks for the medal too :)